# Control of spatiotemporal chaos by stochastic resetting

### v1

@@ Abstract
We study how spatiotemporal chaos in dynamical systems can be controlled by stochastically returning them to their initial conditions. Focusing on discrete nonlinear maps, we analyze how key measures of chaos—the Lyapunov exponent and butterfly velocity, which quantify sensitivity to initial perturbations and the ballistic spread of information, respectively—are reduced by stochastic resetting. We identify a critical resetting rate that induces a dynamical phase transition, characterized by the simultaneous vanishing of the Lyapunov exponent and butterfly velocity, effectively arresting the spread of information. These theoretical predictions are validated and illustrated with numerical simulations of the celebrated logistic map and its lattice extension. Beyond discrete maps, our findings offer insights applicable to a broad class of extended classical interacting systems.
@@
Introduction.– The spatiotemporal scrambling of information in chaotic many-body systems [1, 2, 3, 4, 5] not only underpins the foundations of statistical physics but also poses substantial practical challenges for the smooth operation of information processing devices as their complexity increases [6]. Throughout the history of classical computing, various strategies have been developed to address these challenges and ensure reliable operation within regular dynamical regimes. These include approaches based on feedback and optimal control [7], advanced techniques in control theory [8], stabilizing periodic orbits via probabilistic control [9, 10, 11], and the imposition of dynamical constraints [12, 13]. Although this subject has been extensively studied across diverse communities, out-of-time-order correlations (OTOCs) have recently emerged as a relatively transparent probe of spatiotemporal chaos. They bring the well-developed toolbox of correlation functions to the study of dynamical systems. Initially introduced for quantum mechanical systems in the semiclassical regime [14, 1, 2, 3], OTOCs were subsequently studied in fully quantum systems, such as spin-1/2 chains [15, 16], and more recently applied to purely classical systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. OTOCs quantify the temporal growth and spatial spread of perturbations between two copies of a system prepared with slightly different initial conditions. In classical systems, the exponential sensitivity to initial conditions—a hallmark of chaos—is characterized by Lyapunov exponents, while the ballistic propagation of information is captured by the light-cone structure of OTOCs and quantified by the butterfly velocity. Alongside ongoing efforts to understand the emergence of these two quantities in diverse classical or quantum setups, notably the influence of disorder [27, 28, 29, 30, 31], there is a growing interest in devising strategies to control them through externally engineered protocols.
@@
In this work, we explore how stochastic resetting [32, 33, 34, 35, 36] can be harnessed to delay or suppress the onset of spatiotemporal chaos in classical interacting many-body systems. At its core, stochastic resetting consists in returning the state of a dynamical system to its initial condition at random times. Variations on this protocol include non-Poissonian resetting times, resetting to a cloud of initial conditions, and more. By and large, stochastic resetting has proven to be a versatile nonequilibrium protocol applicable to virtually any dynamical system, with relatively low analytical or computational overhead. It has also been demonstrated experimentally [37, 38, 39].
@@
By introducing a new timescale, stochastic resetting can reshape the system’s transient and steady-state dynamics. In the context of random search problems, it can be used to dramatically reduce the mean first-passage time [40, 41]. From a thermodynamic perspective, stochastic resetting has been articulated within the framework of stochastic thermodynamics [42, 43]. However, much remains to be understood about the impact of stochastic resetting on the dynamics of extended interacting systems [44]. Studies have begun to explore the effects of resetting on correlations and dynamical phase transitions such as phase synchronization in the Kuramoto model [45, 46, 47, 48], Kardar-Parisi-Zhang equation [49], exclusion process [50, 51, 52] and the Ising model [53, 54]. Conversely, stochastic resetting can serve as a resource for dynamically generating correlations between non-interacting particles or fields [55, 56, 57].
@@
We question the fate of chaotic properties in interacting many-body dynamics subject to stochastic resetting, within the framework of discrete maps commonly used in the study of dynamical systems [58, 59, 60]. After formulating the discrete-time dynamics subject to stochastic resetting, we first focus on Lyapunov growth by analyzing zero-dimensional nonlinear maps subject to resetting. Subsequently, we study the spatiotemporal spread of information in extended systems, particularly the butterfly velocity, using coupled map lattices.
@@
Discrete-time maps under stochastic resetting.– Let us consider the dynamics generated by a deterministic discrete-time map interspersed with random resetting events. The resulting stochastic map can be formulated as
$x_{n}\mapsto x_{n+1}=\left\{\begin{array}[]{ll}f(x_{n})&\text{with probability }1-r\\ x_{0}&\text{with probability }r\,,\end{array}\right.$ (3)
@@
for $n\geq 0$. Here, $0\leq r\leq 1$ is the rate at which trajectories are reset to their initial condition $x_{0}$. The state of the system at time $n$ is given by $x_{n}$. Let us first examine how resetting affects the temporal aspect of chaos by considering zero-dimensional maps where $f(x)$ is a non-linear function of $x$. Later, we shall examine the spatial aspects by generalizing to multidimensional-vector states living on a lattice of size $L$, $x_{n}=\{x_{n,\,i}\}_{i=1\ldots L}$. The dynamics in Eq. (3) are explicitly Markovian but this encompasses non-Markovian evolutions with finite time kernel at the cost of adding auxiliary degrees of freedom.
@@
In the presence of resetting ($r>0$), the dynamics in Eq. (3) generate stochastic trajectories. The corresponding ensemble average consists in sampling over the many possible realizations of the resetting times. We find that the stationary distribution is given by [61]
$\tilde{p}_{\rm st}(x)=\sum_{n=0}^{\infty}\tilde{g}_{n}\,\delta(x-x_{n})\,,$ (4)
with $\tilde{g}_{n}:=r(1-r)^{n}\,\Phi(1-r,1,1+n)$ where $\Phi$ is the Hurwitz–Lerch transcendent. Although $\tilde{p}_{\rm st}(x)$ depends on the choice of initial condition $x_{0}$, none of the results presented below depend on $x_{0}$ and we do not perform any average over $x_{0}$. A rigorous assessment of the ergodicity of the resetting dynamics is challenging and, to the best of our knowledge, has not yet been addressed in the literature. On this specific matter, we shall make informed conjectures rather than definitive conclusions. To assess its chaotic behavior, i.e., its sensitivity to perturbations of initial conditions, we define the Lyapunov exponent $\tilde{\lambda}$ as
@@
where $\langle\ldots\rangle_{r}$ denotes the average over resetting times. Unlike the deterministic case, the independence of $\tilde{\lambda}$ with respect to $x_{0}$ is nontrivial and will be clarified below. The computation of $\tilde{\lambda}$ is greatly simplified by expressing $\tilde{d}_{n}:=\langle|{\delta x_{n}}/{\delta x_{0}}|\rangle_{r}$ in Eq. (5) in terms of its deterministic counterpart: $d_{n}:=|{\delta x_{n}}/{\delta x_{0}}|$. This is achieved by the renewal formula
@@
To simplify our analysis, we assume that the deterministic dynamics are characterized by a Lyapunov exponent $\lambda$ of order 1 or greater. This implies that the exponential sensitivity to initial conditions, $|\delta x_{n}/\delta x_{0}|\propto\exp(\lambda n)$, develops as early as $n=1$ and we can neglect the sub-exponential contributions. With this chaotic ansatz in Eq. (6), resetting is found to renormalize the Lyapunov exponent according to
$\tilde{\lambda}=\left\{\begin{array}[]{ll}\lambda+\log(1-r)&\text{ if }r<r_{\rm c}\\ 0&\text{ if }r\geq r_{\rm c}\,,\end{array}\right.$ (9)
where $r_{\rm c}:=1-{\rm e}^{-\lambda}$. This monotonous decay of the Lyapunov with $r$ indicates that increasing the resetting probability systematically reduces the chaoticity of the dynamics. From Eq. (9), it is now clear that $\tilde{\lambda}$ is independent of the initial condition $x_{0}$. $\tilde{\lambda}$ vanishes at a critical resetting rate, $r_{\rm c}\ <1$, signaling a transition from chaotic to non-chaotic dynamics. For $0<r<r_{\rm c}$, the dynamics retain exponential sensitivity to initial conditions, and there remains a finite probability for trajectories to explore the entire phase space, preserving ergodicity. As $r$ approaches $r_{\rm c}$, the Lyapunov vanishes linearly: $\tilde{\lambda}\propto(r-r_{\rm c})^{\mu}$ with the dynamical exponent $\mu=1$. At $r=r_{\rm c}$, a dynamical transition takes place: the exponential sensitivity to initial conditions is lost, and the system shifts to non-chaotic behavior. For $r\geq r_{\rm c}$, the dynamics become non-chaotic and we present arguments in the Supplementary Material [61] indicating a concomitant loss of ergodicity. In this regime, trajectories consistently remain localized near their initial conditions, failing to escape their vicinity.
To numerically test the prediction in Eq. (9), we first reformulate the expression of the Lyapunov exponent in Eq. (5) using the relationship $|\delta x_{n}/\delta x_{0}|=\prod_{p=0}^{n-1}|f^{\prime}(x_{p})|$ for $n\geq 1$. We obtain the explicit expression
@@
This approach only requires analyzing a single trajectory of the non-resetting dynamics, thereby avoiding the complex calculations of linear responses to infinitesimal perturbations and the averaging over numerous resetting trajectories.
Figure 1: Logistic map [Eq. (11), $\alpha=4$] subject to stochastic resetting. (a) Exponential sensitivity to initial perturbations for the deterministic map ($r=0$) and for a finite resetting rate ($r=0.2$). The former validates the exponential ansatz $\log|\delta x_{n}/\delta x_{0}|\propto\lambda\,n$, even at early times. The Lyapunov exponents $\lambda=\log 2$ and $\tilde{\lambda}$ can be extracted from the slopes. (b) The renormalized Lyapunov $\tilde{\lambda}$ computed with Eq. (Control of spatiotemporal chaos by stochastic resetting) is compared to the analytical predictions of Eq. (9) as a function of the resetting rate $r$. The critical resetting is located at $r_{\rm c}=1/2$.
@@
with $x_{0}\in[0,1]$ and $\alpha\in[0,4]$ to ensure the stability of the interval $x\in[0,1]$ under the map. Let us note that the bifurcation diagram of the logistic map when varying $\alpha\in[0,4]$ remains invariant under stochastic resetting, provided an additional average is performed over initial conditions drawn from the stationary distribution of the deterministic map [61]. In practice, we work at $\alpha=4$ where the map exhibits fully developed chaos with a Lyapunov $\lambda=\log 2$ and an ergodic stationary distribution given by $p_{\rm st}(x)={1}/{(\pi\sqrt{x(1-x)})}$ [66]. The results are shown in Fig. 1. The panel (a) validates the ansatz $d_{n}\propto\exp(\lambda\,n)$ that was used to derive the renormalized Lyapunov $\tilde{\lambda}$ in Eq. (9). In the panel (b), we directly compare these analytical predictions to the numerical results, where Lyapunov exponents are computed using Eq. (Control of spatiotemporal chaos by stochastic resetting). The agreement is excellent.
Coupled map lattices (CML).– We now examine the spatial aspect of chaos in the presence of stochastic resetting by considering discrete maps on lattices, commonly referred to as coupled map lattices [67, 68, 69]. For simplicity, we consider one-dimensional lattices of $L$ sites with periodic boundary conditions. The state of the system at time $n$ is now represented by a vector $x_{n}$ with components $x_{n,\,i}$, where $i=1,,\,\ldots,\,L$. We follow both the temporal growth and the spatial spread of a perturbation originating from site $j$ by means of the classical OTOCs defined as
@@
for $n\geq|i-j|$, and $D_{n,\,ij}=0$ otherwise. Here, the function $\lambda(v)$ is commonly referred to as a velocity-dependent Lyapunov exponent. This ansatz is known to capture the phenomenology in a wide class of models [70, 71, 72, 23]. We assume that $\lambda(v)$ is a continuous monotonically decreasing function with negative curvature, $\lambda(0)>0$, and that $\lambda(v)$ crosses zero at a finite velocity $v_{\rm B}$. At large times, this ansatz describes a traveling wavefront that propagates ballistically at the so-called butterfly velocity $v_{\rm B}=\lambda^{-1}(0)$, where $\lambda^{-1}$ denotes the inverse function of $\lambda$. At this front, the exponential growth of the OTOC is characterized by the Lyapunov rate $\lambda_{\rm B}=-v_{\rm B}\lambda^{\prime}(v_{\rm B})>0$ [61]. Similarly to Eq. (6), the OTOC satisfies the following renewal equation for $n\geq 1$
@@
$\tilde{v}_{\rm B}=\left\{\begin{array}[]{ll}\lambda^{-1}\left(-\log(1-r)\right)&\text{ if }r<r_{\rm c}\\ 0&\text{ if }r\geq r_{\rm c}\,,\end{array}\right.$ (17)
@@
In the phase $r<r_{\rm c}$, both the butterfly velocity and the Lyapunov are reduced by stochastic resetting. The near-critical behavior is governed by the behavior of $\lambda(v)$ in the vicinity of $v=0$. Assuming it is described by $\lambda(v)-\lambda(0)\sim v^{\alpha}$ where $\alpha\geq 1$, the renormalized butterfly velocity and Lyapunov exponent scale as $\tilde{v}_{\rm B}\sim(r_{\rm c}-r)^{\alpha}$ and $\tilde{\lambda}_{\rm B}\sim(r_{\rm c}-r)^{\alpha^{2}}$. In the phase $r>r_{\rm c}$, the resetting rate becomes sufficiently strong to drive a dynamical phase transition where the dynamics become non-chaotic, non-ergodic and the spread of information is brought to a complete arrest. Explicit computations and additional details are provided in the Supplementary Material [61].
Figure 2: Coupled logistic map [Eq. (20] subject to stochastic resetting: the OTOC $\tilde{D}_{n,\,ij}$ as a function of space and time reveals the ballistic spreading of perturbations. (a) Deterministic case, $r=0$. (b) Finite resetting rate $r=0.2<r_{\rm c}\approx 0.38$. (c) $r=0.4>r_{\rm c}$. The butterfly velocities can be extracted from the slopes of the light cones. Lattice size $L=701$, $\alpha=4$, and $c=0.1$. The data is averaged over the initial perturbation site $j$, ranging from 1 to $L$.
@@
with periodic boundary conditions. The first onsite term is precisely the nonlinearity in Eq. (11). The second term can be seen as a discrete Laplace operator with the diffusive coupling $c\in[0,1]$. We use a single initial state with random $x_{0,\,i}\in[0,1]$ for $i=1,\ldots,L$. The map preserves $x_{n,\,i}\in[0,1]$ at all times. The coupled logistic map exhibits a rich phase diagram and has been widely studied in the context of spatiotemporal chaos, addressing phenomena such as spatial bifurcation, pattern selection, spatiotemporal intermittency, and soliton turbulence [67, 68, 69, 73, 74, 75]. Here, we fix $\alpha=4$ and $c=0.1$, corresponding to a regime of fully developed turbulence. Additional details are provided in the supplementary material [61]. In the presence of resetting, the OTOCs are numerically obtained with the renewal formula in Eq. (14), making use of the OTOCs of the deterministic map. The latter are computed using an algorithm discussed in the Supplementary Material [61]. To avoid boundary effects, we restrict times to $2n<L$.
Figure 2 illustrates the spatiotemporal behavior of the OTOC $\tilde{D}_{n,\,ij}$ under different resetting conditions. Panel (a) shows the light-cone structure of the OTOC in the absence of resetting ($r=0$). The local perturbation of the initial condition evolves into a traveling wavefront that propagates ballistically at the butterfly velocity $v_{\rm B}$ visible as the slope of the light cone at late times. Note that the causal light cone at the maximal Lieb-Robinson velocity $v_{LR}=1$ can be seen at early times, and it is overshadowed by the prominence of the butterfly front at later times. In Panel (b), the butterfly velocity is visibly reduced by a finite resetting rate ($r=0.2$). In Panel (c), for $r>r_{\rm c}$, the light cone collapses as the dynamics freeze, signaling the arrest of spatiotemporal chaos. Noteworthy, a similar dynamical phase transition was reported in deterministic classical many-body systems with kinetic constraints [12, 13].
Figure 3: Spatiotemporal chaos in the logistic map lattice subject to stochastic resetting. (a) The velocity-dependent Lyapunov $\lambda(v)$ [Eq. (13)] is numerically extracted from the collapse of $(\log D_{n,\,ij})/n$ versus $v:=|i-j|/n$ for various values of $i-j$ in the deterministic case ($r=0$). The butterfly velocity $\tilde{v}_{\rm B}$ at $r>0$ can be determined by solving $\lambda(\tilde{v}_{\rm B})=-\log(1-r)$. The slope of $\lambda(v)$ at this solution determines the Lyapunov rate $\tilde{\lambda}_{\rm B}$. (b) $\tilde{v}_{\rm B}$ is extracted from the slope of the light cones (see Fig. 2) and is compared to the prediction of Eq. (17) as a function of the resetting rate $r$. Lattice size $L=701$, $\alpha=4$, and $c=0.1$.

### v2

@@ Abstract
We study how spatiotemporal chaos in dynamical systems can be controlled by stochastically returning them to their initial conditions. Focusing on discrete nonlinear maps, we analyze how key measures of chaos—the Lyapunov exponent and butterfly velocity, which quantify sensitivity to initial perturbations and the ballistic spread of information, respectively—are reduced by stochastic resetting. We identify a critical resetting rate that induces a dynamical phase transition, characterized by the simultaneous vanishing of the Lyapunov exponent and butterfly velocity, effectively arresting the spread of information. These theoretical predictions are validated and illustrated with numerical simulations of the celebrated logistic map and its lattice extension. Beyond discrete maps, our findings are applicable to virtually any chaotic extended classical many-body system.
I Introduction
@@
The spatiotemporal scrambling of information in chaotic many-body systems [1, 2, 3, 4, 5] not only underpins the foundations of statistical physics but also poses substantial practical challenges for the smooth operation of information processing devices as their complexity increases [6]. Throughout the history of classical computing, various strategies have been developed to address these challenges and ensure reliable operation within regular dynamical regimes. These include approaches based on feedback and optimal control [7, 8, 9], advanced techniques in control theory [10], stabilizing periodic orbits via probabilistic control [11, 12, 13], and the imposition of dynamical constraints [14, 15]. Although this subject has been extensively studied across diverse communities, out-of-time-order correlations (OTOCs) have recently emerged as a relatively transparent probe of spatiotemporal chaos. They bring the well-developed toolbox of correlation functions to the study of dynamical systems. Initially introduced for quantum mechanical systems in the semiclassical regime [16, 1, 2, 3], OTOCs were subsequently studied in fully quantum systems, such as spin-1/2 chains [17, 18], and more recently applied to purely classical systems [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. OTOCs quantify the temporal growth and spatial spread of perturbations between two copies of a system prepared with slightly different initial conditions. In classical systems, the exponential sensitivity to initial conditions—a hallmark of chaos—is characterized by Lyapunov exponents, while the ballistic propagation of information is captured by the light-cone structure of OTOCs and quantified by the butterfly velocity. Alongside ongoing efforts to understand the emergence of these two quantities in diverse classical or quantum setup [29, 30, 31, 32, 33], there is a growing interest in devising strategies to control them through externally engineered protocols.
@@
In this work, we explore how stochastic resetting [34, 35, 36, 37, 38, 39] can be harnessed to delay or suppress the onset of spatiotemporal chaos in classical interacting many-body systems. At its core, stochastic resetting consists in returning the state of a dynamical system to its initial condition at random times. Variations on this protocol include non-Poissonian resetting times, resetting to a cloud of initial conditions, and more. Stochastic resetting can also arise naturally in the effective description of deterministic chaotic maps [40, 41]. By and large, stochastic resetting has proven to be a versatile nonequilibrium protocol applicable to virtually any dynamical system, with relatively low analytical or computational overhead. It has also been demonstrated experimentally [42, 43, 44].
@@
By introducing a new timescale, stochastic resetting can reshape the system’s transient and steady-state dynamics. In the context of random search problems, it can be used to dramatically reduce the mean first-passage time [45, 46]. From a thermodynamic perspective, stochastic resetting has been articulated within the framework of stochastic thermodynamics [47, 48]. However, much remains to be understood about the impact of stochastic resetting on the dynamics of extended interacting systems [49]. Studies have begun to explore the effects of resetting on correlations and dynamical phase transitions such as phase synchronization in the Kuramoto model [50, 51, 52], Kardar-Parisi-Zhang equation [53], exclusion process [54, 55, 56] and the Ising model [57, 58]. Conversely, stochastic resetting can serve as a resource for dynamically generating correlations between non-interacting particles or fields [59, 60, 61].
@@
We question the fate of chaotic properties in interacting many-body dynamics subject to stochastic resetting, within the framework of discrete maps commonly used in the study of dynamical systems [62, 63, 64]. After formulating the discrete-time dynamics subject to stochastic resetting [37], we first focus on Lyapunov growth by analyzing zero-dimensional nonlinear maps subject to resetting. Subsequently, we study the spatiotemporal spread of information in extended systems, particularly the butterfly velocity, using coupled map lattices.
II Discrete-time maps under stochastic resetting
@@
Let us consider the dynamics generated by a deterministic discrete-time map interspersed with random resetting events. The resulting stochastic map can be formulated as
[…]
@@
where $x_{n}$ denotes the state of the system at time $n\geq 0$. Here, $0\leq r\leq 1$ is the probability, at each time step, to reset the system to its initial condition $x_{0}$. Let us first examine how resetting affects the temporal aspect of chaos by considering zero-dimensional maps where $f(x)$ is a non-linear function of the state $x$. Later, we shall examine the spatial aspects by generalizing to multidimensional-vector states living on a lattice of size $L$, $x_{n}=\{x_{n,\,i}\}_{i=1\ldots L}$. The dynamics in Eq. (3) are explicitly Markovian but this encompasses non-Markovian evolutions with finite time kernel at the cost of adding auxiliary degrees of freedom.
@@
In the presence of resetting ($r>0$), the dynamics in Eq. (3) generate stochastic trajectories. The corresponding ensemble average consists in sampling over the many possible realizations of the resetting times. We show in Appendix A.2 that the stationary distribution is given by
$\tilde{p}_{\rm st}(x)=\sum_{n=0}^{\infty}\tilde{g}_{n}\,\delta(x-f^{n}(x_{0}))\,$, (4)
with $\tilde{g}_{n}:=r(1-r)^{n}$ and $f^{n}(x_{0})$ stands for the $n$-th iteration of the deterministic map $f(f(\ldots f(x_{0})))$. Although $\tilde{p}_{\rm st}(x)$ depends on the choice of initial condition $x_{0}$, none of the results presented below depend on $x_{0}$ and we do not perform any average over $x_{0}$. A rigorous assessment of the ergodicity of the resetting dynamics is challenging and, to the best of our knowledge, has not yet been addressed in the literature [65, 66, 67, 68]. On this specific matter, we shall make informed conjectures rather than definitive conclusions. To assess its chaotic behavior, i.e., its sensitivity to perturbations of initial conditions, we define the Lyapunov exponent $\tilde{\lambda}$ as
@@
where $\langle\ldots\rangle_{r}$ denotes the average over resetting times. Note that, since the dynamics intrinsically comprises both deterministic evolution and stochastic events, the averaging is carried out before taking the logarithm. Unlike the deterministic case, the independence of $\tilde{\lambda}$ with respect to $x_{0}$ is not trivial and will be clarified below. The computation of $\tilde{\lambda}$ is greatly simplified by expressing $\tilde{d}_{n}:=\langle|{\delta x_{n}}/{\delta x_{0}}|\rangle_{r}$ in Eq. (5) in terms of its deterministic counterpart: $d_{n}:=|{\delta x_{n}}/{\delta x_{0}}|$. This is achieved by the renewal formula
@@
To simplify the derivation of explicit results, we assume that the deterministic dynamics are characterized by a Lyapunov exponent $\lambda$ of order one or larger. Under this assumption, the exponential sensitivity to initial conditions, $|\delta x_{n}/\delta x_{0}|\propto\exp(\lambda n)$, develops as early as $n=1$ and we can neglect the sub-exponential contributions. It should be emphasized, however, that the validity of the results is broader, and the restriction to large $\lambda$ will be relaxed later. With this chaotic ansatz in Eq. (6), resetting is found to renormalize the Lyapunov exponent according to
[…]
where $r_{\rm c}:=1-{\rm e}^{-\lambda}$. This monotonous decay of the Lyapunov with $r$ indicates that increasing the resetting probability systematically reduces the chaoticity of the dynamics. From Eq. (9), it is now clear that $\tilde{\lambda}$ is independent of the initial condition $x_{0}$. $\tilde{\lambda}$ vanishes at a critical resetting rate, $r_{\rm c}\ <1$, signaling a transition from chaotic to non-chaotic dynamics. For $0<r<r_{\rm c}$, the dynamics retain exponential sensitivity to initial conditions, and there remains a finite probability for trajectories to explore the entire phase space, preserving ergodicity. As $r$ approaches $r_{\rm c}$, the Lyapunov vanishes linearly: $\tilde{\lambda}\propto(r_{\rm c}-r)^{\mu}$ with the dynamical exponent $\mu=1$. At $r=r_{\rm c}$, a dynamical transition takes place: the exponential sensitivity to initial conditions is lost, and the system shifts to non-chaotic behavior. For $r\geq r_{\rm c}$, the dynamics become non-chaotic and we present arguments in Appendix A.3 indicating a concomitant loss of ergodicity. In this regime of vanishing Lyapunov, unlike the standard periodic attractor or fixed point pictures, the trajectories are stochastic and remain confined to sequences that revisit the initial state and its early iteration states. Although generically, a Lyapunov exponent can be negative, Eq. (9) predicts non-negative values. This is because resetting two copies of the map to their respective infinitesimally close initial conditions hinders trajectories from approaching each other exponentially in time.
To numerically test the prediction in Eq. (9), we first reformulate the expression of the Lyapunov exponent in Eq. (5) using the chain rule $|\delta x_{n}/\delta x_{0}|=\prod_{p=0}^{n-1}|f^{\prime}(x_{p})|$ for $n\geq 1$. We obtain an explicit expression that does not rely on any specific assumption regarding $\lambda$,
@@
This approach only requires analyzing a single trajectory of the non-resetting dynamics, thereby avoiding the complex calculations of linear responses to infinitesimal perturbations and the averaging over numerous resetting trajectories. Equation (II) reveals that $\tilde{\lambda}$ is dominated by those trajectories with a small number of resetting events. This can be seen by inspecting the contributions of the last two terms after substituting into the equation the chaotic ansatz at large $n$, $d_{n}\propto\exp(\lambda n)$, and using that $d_{n}$ is at most exponential at early and intermediate times.
Figure 1: Logistic map [Eq. (11), $\alpha=4$] subject to stochastic resetting. (a) Exponential sensitivity to initial perturbations for the deterministic map ($r=0$) and for a resetting rate $r=0.2<r_{\rm c}$. The former validates the exponential ansatz $d_{n}:=\log|\delta x_{n}/\delta x_{0}|\propto\lambda\,n$, even at early times. The Lyapunov exponents $\lambda=\log 2$ and $\tilde{\lambda}$ can be extracted from the slopes. For $r=0.6>r_{\rm c}$, $\tilde{d}_{n}$ saturates to a constant. (b) The renormalized Lyapunov $\tilde{\lambda}$ computed with Eq. (II) is compared to the analytical predictions of Eq. (9) as a function of the resetting rate $r$. The critical resetting is located at $r_{\rm c}=1/2$.
@@
with the initial state $x_{0}\in[0,1]$ and $\alpha\in[0,4]$ to ensure the stability of the interval $x\in[0,1]$ under the map. Let us note that the bifurcation diagram of the logistic map when varying $\alpha\in[0,4]$ remains invariant under stochastic resetting, provided an additional average is performed over initial conditions drawn from the stationary distribution of the deterministic map, see Appendix A.2. In practice, we work at $\alpha=4$ where the map exhibits fully developed chaos with a Lyapunov $\lambda=\log 2$ and an ergodic stationary distribution given by $p_{\rm st}(x)={1}/{(\pi\sqrt{x(1-x)})}$ [73]. We emphasize that the analytics leading to Eq. (9) holds when the bare Lyapunov exponent (i.e. in the absence of resetting) is positive, and any value of $\alpha$ satisfying this condition will guarantee similar results. The results are shown in Fig. 1. Panel (a) validates the ansatz $d_{n}\propto\exp(\lambda\,n)$ that was used to derive the renormalized Lyapunov $\tilde{\lambda}$ in Eq. (9). In Panel (b), we directly compare these analytical predictions to the numerical results, where Lyapunov exponents are computed using Eq. (II). The agreement is excellent. When $r>r_{\rm c}$, the loss of ergodicity is signaled by $\tilde{d}_{n}$ plateauing to a finite value which can be interpreted as a quantifier of the volume fraction of the effectively accessible phase space.
III Coupled map lattices
We now examine the spatial aspect of chaos in the presence of stochastic resetting by considering discrete maps on lattices, commonly referred to as coupled map lattices [74, 75, 76]. For simplicity, we consider one-dimensional lattices of $L$ sites with periodic boundary conditions. The state of the system at time $n$ is now represented by a vector $x_{n}$ with components $x_{n,\,i}$, where $i=1,,\,\ldots,\,L$. We follow both the temporal growth and the spatial spread of a perturbation originating from site $j$ by means of the classical OTOCs defined as
@@ III Coupled map lattices
for $n\geq|i-j|$, and $D_{n,\,ij}=0$ otherwise. Here, the function $\lambda(v)$ is commonly referred to as a velocity-dependent Lyapunov exponent. This ansatz is known to capture the phenomenology in a wide class of models [77, 78, 26, 25]. We assume that $\lambda(v)$ is a continuous monotonically decreasing function with negative curvature, $\lambda(0)>0$, and $\lambda(v)$ crosses zero at a finite velocity $v_{\rm B}$. At large times, this ansatz describes a traveling wavefront that propagates ballistically at the so-called butterfly velocity $v_{\rm B}=\lambda^{-1}(0)$, where $\lambda^{-1}$ denotes the inverse function of $\lambda$. At this front, the exponential growth of the OTOC is characterized by the Lyapunov rate $\lambda_{\rm B}=-v_{\rm B}\lambda^{\prime}(v_{\rm B})>0$, see the details in Appendix C.2. Similarly to Eq. (6), the OTOC satisfies the following renewal equation for $n\geq 1$
@@
[…]
@@
In the phase $r<r_{\rm c}$, both the butterfly velocity and the Lyapunov are reduced by stochastic resetting. The near-critical behavior is governed by the behavior of $\lambda(v)$ in the vicinity of $v=0$. Assuming it is described by $\lambda(v)-\lambda(0)\sim v^{\alpha}$ where $\alpha\geq 1$, the renormalized butterfly velocity and Lyapunov exponent scale as $\tilde{v}_{\rm B}\sim(r_{\rm c}-r)^{1/\alpha}$ and $\tilde{\lambda}_{\rm B}\sim(r_{\rm c}-r)^{\mu}$, with $\mu=1$. In the phase $r>r_{\rm c}$, the resetting rate becomes sufficiently strong to drive a dynamical phase transition where the dynamics become non-chaotic, non-ergodic and the spread of information is brought to a complete arrest. Explicit computations and additional details are provided in the Appendix C.2.
Figure 2: Coupled logistic map [Eq. (20] subject to stochastic resetting: the OTOC $\tilde{D}_{n,\,ij}$ as a function of space and time reveals the ballistic spreading of perturbations for the (a) Deterministic case at $r=0$, and (b) Finite resetting rate $r=0.2<r_{\rm c}\approx 0.38$. (c) Dynamical arrest at $r=0.4>r_{\rm c}$. The butterfly velocities can be extracted from the slopes of the light cones. Lattice size $L=701$, $\alpha=4$, and $c=0.1$. The data are spatially averaged over the initial perturbation site $j$, ranging from 1 to $L$.
@@
with periodic boundary conditions. The first onsite term is precisely the nonlinearity in Eq. (11). The second term can be seen as a discrete Laplace operator with the diffusive coupling $c\in[0,1]$. We use a single initial state with random $x_{0,\,i}\in[0,1]$ for $i=1,\ldots,L$. The map preserves $x_{n,\,i}\in[0,1]$ at all times. The coupled logistic map exhibits a rich phase diagram and has been widely studied in the context of spatiotemporal chaos, addressing phenomena such as spatial bifurcation, pattern selection, spatiotemporal intermittency, and soliton turbulence [74, 75, 76, 79, 80, 81, 82]. Here, we fix $\alpha=4$ and $c=0.1$, corresponding to a regime of fully developed turbulence. Additional details are provided in the Appendix B. In the presence of resetting, the OTOCs are numerically obtained with the renewal formula in Eq. (14), making use of the OTOCs of the deterministic map. The latter are computed using an algorithm discussed in the Appendix B. We restrict times to $2\tilde{v}_{{\rm B}}n<L$ to exclude boundary effects.
Figure 3: Spatiotemporal chaos in the coupled logistic map subject to stochastic resetting. (a) The velocity-dependent Lyapunov $\lambda(v)$ [Eq. (13)] is numerically extracted from the collapse of $(\log D_{n,\,ij})/n$ versus v:=|ij|/nv:=|i-j|/n for various values of iji-j in the deterministic case ($r=0$). The butterfly velocity $\tilde{v}_{\rm B}$ at $r>0$ can be determined by solving $\lambda(\tilde{v}_{\rm B})=-\log(1-r)$. The slope of $\lambda(v)$ at this solution determines the Lyapunov rate $\tilde{\lambda}_{\rm B}$. (b) $\tilde{v}_{\rm B}$ is extracted from the slope of the light cones (see Fig. 2) and is compared to the prediction of Eq. (17) as a function of the resetting rate $r$. Lattice size $L=701$, $\alpha=4$, and $c=0.1$.
Figure 2 illustrates the spatiotemporal behavior of the OTOC $\tilde{D}_{n,\,ij}$ under different resetting conditions. Panel (a) shows the light-cone structure of the OTOC in the absence of resetting ($r=0$). The local perturbation of the initial condition evolves into a traveling wavefront that propagates ballistically at the butterfly velocity $v_{\rm B}$ visible as the slope of the light cone at late times. Note that the causal light cone at the maximal Lieb-Robinson velocity $v_{\rm LR}=1$ can be seen at early times, and it is overshadowed by the prominence of the butterfly front at later times. In Panel (b), the butterfly velocity is visibly reduced by a finite resetting rate ($r=0.2$). In Panel (c), for $r>r_{\rm c}$, the OTOC quickly saturates to a function $\tilde{D}_{\infty,\,ij}:=\lim_{n\to\infty}\tilde{D}_{n,\,ij}$ that is exponentially localized around the perturbation site $j$. The light cone collapses as the dynamics freeze, signaling the arrest of spatiotemporal chaos. Noteworthy, a similar dynamical phase transition was reported in deterministic classical many-body systems with kinetic constraints [14, 15].